20g : 100g = 1metre : 500cm
 20g : 100g = 1metre : 500cm

Grade:12

1 Answers

Harshit Singh
2 years ago
Dear Student

The given statement is True
Consider the given ratios, 20g : 100g = 1metre : 500cm

20/100 = 1/500 We know that, 1 metre = 100 cm

So, 20g: 100g = 100cm : 500 cm
20/100 = 100/500
20/100 is further simplified by dividing both numerator and denominator by 20 we get, = 1/5
100/500 is further simplified by dividing both numerator and denominator by 100 we get,
= 1/5
Hence, 1/5 = 1/5
Therefore, 20g : 100g = 1metre : 500cm

Thanks